OpenStudy (anonymous):

Explain, in complete sentences, how you would use the elimination method to solve the following system of equations. Provide the solution to the system. 5x – 3y = 4 3x – 5y = –4

OpenStudy (anonymous):

will someone please help

OpenStudy (anonymous):

Alright. I'll show you the steps, but you're gonna have to interpret my steps and write sentences for it. 5x – 3y = 4 3x – 5y = –4 Multiply the equations by something so that when adding them, a variable cancels out. -3(5x – 3y = 4) 5(3x – 5y = –4) -15x + 9y = -12 15x - 25y = -20 Add vertically. 0 - 16y = -32 THis is the identity property of zero. -16y = -32 Divide by -16 to get y isolated. y = 2 Plug this into the equation and solve for y to solve for x. 5x – 3y = 4 5x – 3(2) = 4 5x - 6 = 4 5x = 10 x = 2 Answers: x = 2, y = 2 Solution Set: (2, 2)